OpenStudy (anonymous):

The height of the Empire State Building is 318.00 meters. If a stone is dropped from the top of the building, what is the stone's velocity just before it strikes the ground?

OpenStudy (shane_b):

Since this is a constant acceleration you can use the kinematic equation:\[V_f^2=V_i^2+2ad\]\[V_f^2=(0m/s)^2+2(9.8m/s^2)(318.00m)\]\[V_f^2=2(9.8m/s^2)(318.00m)\]\[V_f=\sqrt{2(9.8m/s^2)(318.00m)}\]